# Diffusion Models in Simulation-Based Inference: A Tutorial Review

### v1

@@ 4.3.1 Multiscale Estimation
Many simulation-based inference problems involve high-dimensional, spatially structured parameters, where the resolution can be modeled hierarchically. So instead of multiple observations, we have one high-dimensional observation $\mathbf{y}_{\text{obs}}$, which can be mapped to different resolutions using a downscaling function $h(\mathbf{y}_{\text{obs}},r)$. We can then offload generation across multiple scales: we start at a coarse resolution and progressively upscale while refining detail. This idea was popularized in conditional image generation by cascaded diffusion pipelines that train separate (class-conditional) models at $(64\!\to\!128\!\to\!256)\,\text{px}$ resolution stages, conditioning each stage on the upscaled output of the previous one [Ho2021-cascaded].
In the SBI setting, [cirakman2025efficient] combine diffusion models with wavelet decompositions to infer high-resolution velocity fields from seismic data. Here a diffusion model is learned for the top-level coarse component $p(\boldsymbol{\eta}^{(R)}\mid h(\mathbf{y}_{\text{obs}},R))$ and, for each resolution level $r$, a conditional detail model $p(\boldsymbol{\theta}^{(r)}\mid\boldsymbol{\eta}^{(r)},h(\mathbf{y}_{\text{obs}},r))$. This can be combined with an inverse wavelet transform to obtain the conditioning variable for the next finer stage $\operatorname{WT}^{-1}(\boldsymbol{\eta}^{(r)},\boldsymbol{\theta}^{(r)})=\boldsymbol{\eta}^{(r-1)}$. The joint posterior over all resolution levels factorizes as
$p(\{\boldsymbol{\eta}^{(r)}\}_{r=0}^{R},\{\boldsymbol{\theta}^{(r)}\}_{r=1}^{R}\mid\mathbf{y}_{\text{obs}})$ $\;=\;p(\boldsymbol{\eta}^{(R)}\mid h(\mathbf{y}_{\text{obs}},R))\;p(\boldsymbol{\theta}^{(R)}\mid\boldsymbol{\eta}^{(R)},h(\mathbf{y}_{\text{obs}},R))$
$\qquad\times\prod_{r=1}^{R-1}p(\boldsymbol{\theta}^{(R-r)}\mid\boldsymbol{\eta}^{(R-r)},h(\mathbf{y}_{\text{obs}},R-r)),$ (40)
so that training and ancestral sampling proceed naturally from coarse to fine scales. Each score estimator can be trained separately due to the exact factorization above and the analytic deconstruction path from high-resolution training data via $\operatorname{WT}(\boldsymbol{\eta}^{(r)})=(\boldsymbol{\eta}^{(r+1)},\boldsymbol{\theta}^{(r+1)})$. While the high-dimensional posterior has long-range spatial correlations, the individual conditional posteriors are closer to Gaussian distributions even when the high-dimensional posterior is not, making estimation of the individual factors easier [Guth2022-waveletScore-basedGenerativeModelingv]. This idea was first demonstrated with normalizing flows by [yu2020wavelet], and [Guth2022-waveletScore-basedGenerativeModelingv] extended it to score-based generative models. Diffusion models are particularly well suited for this setting, as they scale to high dimensions and are very flexible in how the conditioning is used within the neural network architecture.
@@ 4.3.2 Causal Posterior Estimation
Causal posterior estimation extends simulation-based inference to settings where the posterior distribution inherits the structure of a causal model [dirmeier2025causal]. A key idea is to integrate the causal directed acyclic graph (DAG) directly into the network architecture of the diffusion model and the probability flow.
@@ 4.3.3 Joint Estimation
Instead of training only a conditional model for $p(\boldsymbol{\theta}\mid\mathbf{y})$, one can train a diffusion model on the joint $\mathbf{z}=(\boldsymbol{\theta},\mathbf{y})$ [gloeckler2024all]. Using a transformer as the neural network architecture for the diffusion model, one can process any subset of factors forming the joint distribution. Yet, this architectural choice is prohibitive for constrained architectures (e.g., normalizing flows), as they typically require explicit conditioning on fixed inputs rather than being able to marginalize variables implicitly. In practice, the network is trained with random masking over components of $\mathbf{z}^{M_{C}}_{t}=(1-M_{C})\cdot\mathbf{z}_{t}+M_{C}\cdot\mathbf{z}_{0}$, where $M_{C}\in\{0,1\}^{d}$, leading to the modified objective
$\hat{s}=\operatorname*{argmin}_{s}\,\mathbb{E}_{M_{C}\sim p(M_{C}),\,\mathbf{z}_{0}\sim p(\mathbf{z}_{0}),\,t\sim\mathcal{U}(0,1),\,\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})}\left[\,\omega_{t}\,\|(1-M_{C})\cdot\bigl(s(\mathbf{z}^{M_{C}}_{t},t)-\nabla_{\mathbf{z}_{t}}\!\log p(\mathbf{z}_{t}\mid\mathbf{z}_{0})\bigr)\|_{2}^{2}\,\right].$ (43)
@@
Beyond posterior approximation, joint training also enables Bayesian model comparison using both the likelihood and the posterior [gunes2025compass]. Having access to all conditional distributions makes this approach particularly appealing for sequential neural score estimation. For example, [sharrock2024sequential] introduced SNPSE-C, which incorporates a score-based correction into the denoising score matching objective to refine posterior estimates iteratively. SNPSE-C requires access to likelihood and prior scores, which introduces additional simplifications and approximations, but these can be avoided when the diffusion model is trained directly on the joint. Extensions to spatial and non-parametric inference have also been proposed. For instance, [Tesso2025-spatformerSimulation-basedInferenceTransformersSpatialStatisticsi] adapted the framework to spatially correlated data and demonstrated inference over multiple spatial conditional distributions simultaneously by providing temporal and spatial information as inputs to a transformer-based summary network.
@@ 4.3.4 Learning Scores Beyond Euclidean Spaces
The concept of the score function can be generalized beyond Euclidean parameter spaces to distributions defined on other structured domains. A detailed survey of diffusion models for data with special structures unrelated to SBI is provided by [yang2023diffusion], covering cases such as discrete variables, symmetry-invariant data, and data constrained to manifolds. For targets with intrinsic geometric constraints, the score can be formulated directly on a manifold equipped with its metric structure. Examples include $\operatorname{E}(3)$-equivariant diffusion models [hoogeboom2022equivariant], which respect rigid-body symmetries, and Riemannian score-based models [de2022riemannian], where both forward and reverse processes are defined with respect to the manifold geometry. [jagvaral2022modeling] demonstrated this approach for the estimation of astrophysical orientation distributions on $\operatorname{SO}(3)$ extendable to neural likelihood estimation (NLE).
The score can also be defined for infinite-dimensional spaces when the target variable is a function rather than a finite-dimensional vector. [Pidstrigach2023-infinite-dimensionalDiffusionModelsq] introduced diffusion models for Hilbert spaces and applied them to generative modeling of functions for Bayesian inverse problems and SBI. This formulation enables posterior estimation for high-dimensional objects like spatial fields, time series, or solutions to partial differential equations. A notable subtlety is that in infinite dimensions, the score matching objective is no longer equivalent to the standard denoising score matching loss, which requires careful reformulation of the training objective.
To address the fragmentation of generative objectives across different structured domains, Generator Matching reformulates the learning problem as matching the infinitesimal generator of a Markov process [holderrieth2024generator]. By targeting the linear operator governing the stochastic evolution rather than specific probability paths or noise distributions, this framework provides a unified objective for neural likelihood estimation on arbitrary state spaces, in particular it allows the construction of multi-modal generative models.
@@ 5 Design Choices for Diffusion Models in Simulation-Based Inference
An overarching decision concerns the inference backbone, namely the network responsible for sampling the target quantities $\mathbf{z}\sim p(\mathbf{z}\mid\mathbf{x})$. In principle, any architecture capable of handling the structure of the condition $\mathbf{x}$ is admissible, for instance, hybrid convolutional-recurrent models for time series [zhang2023solving], or permutation-invariant networks for sets [zaheer2017deep, lee2019set]. A common strategy for posterior estimation is to separate the model into a summary network and an inference network. The summary network compresses the data into a fixed-dimensional representation (i.e., an encoder or an embedding network), and the inference network is typically implemented as a stack of (residual) MLPs acting on this summary [radev2020bayesflow, chen2023learning]. However, this separation is optional; the flexibility of diffusion models allows these components to be folded into a single transformer architecture [gloeckler2024all].
@@
Figure 7: Designing diffusion models for simulation-based inference (SBI). A Variance-preserving (VP) EDM schedule. B Variance-exploding (VE) EDM schedule (until $\sigma_{1}=131.9$). C Noise schedules of the log signal-to-noise ratio (SNR). The cosine and flow matching schedules exhibit similar relationships between time and log SNR, yet differ markedly in how the log SNR is weighted in the loss. D Flow matching emphasizes regions of low SNR more strongly than the other schedules. This can be adapted by sampling time $t$ from a power-law distribution with $\rho=-0.6$ rather than uniformly, which aligns the effective weighting with that of the EDM schedule. E Deterministic and stochastic latent trajectories from noise ($t=1$) to the posterior ($t=0$) for diffusion models trained with different noise schedules. The learned trajectories also reflect these design differences. Flow matching with optimal transport (OT) produces relatively straight paths of the probability flow from noise to the posterior, while EDM VP tends to give straighter paths than EDM VE. Stochastic sampling displays highly irregular motion. The example problem shown here is the conditional two moons [lueckmann2021benchmarking] and the stochastic trajectories are smoothed with a simple moving average.
@@ 5.1 Variance of the Noise Schedule
A core design choice in diffusion models concerns the denoising process itself. The forward process (5) is characterized by the functions $\alpha_{t}$ and $\sigma_{t}$, which together specify the noise schedule. These functions determine how signal and noise are blended over time, and thereby shape both the difficulty of the denoising task and the model’s overall behavior [SongSoh2020]. In the variance-preserving (VP) parameterization,
@@
the total variance $\alpha_{t}^{2}+\sigma_{t}^{2}$ remains constant across $t$, assuming $\operatorname{Var}(\mathbf{z}_{0})=1$ and $0\leq\sigma_{t}\leq 1$ (Figure 7A). This ensures a balance between signal and noise at every step and underlies the original diffusion model formulation [HoJai2020a], which is still widely used in practice.
@@
so that the total variance grows throughout the forward process (Figure 7B). In some low-dimensional SBI inference tasks, VE has been reported to outperform VP [sharrock2024sequential, gloeckler2024all], while for high-dimensional imaging data, VP schedules generally perform better than VE schedules [SongSoh2020]. In our experiments Section 6.1, we find the VP schedule to be superior in most SBI tasks.
@@
which yields a lower total variance than the VP case for all $t$. For density estimation tasks, [SongSoh2020] observed that a sub-VP noise schedule can improve accuracy. However, most current SBI papers adopt either a VP schedule combined with a linear noise schedule or a VE schedule, and only a single paper employed a sub-VP schedule [Zeng2025-amortizedLikelihood-freeStochasticModelUpdatingViaConditionalDiffusionModelStructuralHealthMonitoringw].
In practice, the choice of variance type interacts with both the noise schedule and the weighting function, jointly shaping the training difficulty over $t$, as we will discuss in the following sections. In our case studies, we find that for most problems, the VP schedule is a suitable choice (see Section 6). [lu2025simplifying] show that the variance type does not fundamentally affect sample quality, since any formulation can be transformed into a VP schedule by rescaling $\alpha_{t}$ and $\sigma_{t}$ with $\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}$ and adjusting the weighting function accordingly. They introduce a trigonometric reparameterization, TrigFlow, which maps arbitrary schedules into VP form. Moreover, [tang2024contractive] proposed contractive variants of VP and sub-VP schedules that suppress the propagation of score and discretization errors, albeit at the cost of introducing possible bias.
@@ 5.2 Noise Schedule and Signal-to-Noise Ratio
The noise schedule determines how the log signal-to-noise ratio (SNR), $\lambda_{t}=\log(\alpha_{t}^{2}/\sigma_{t}^{2})$, evolves over time $t$, and thus controls the relative difficulty of denoising at different steps. Over the past few years, several schedules have been proposed, each with different motivations and practical tradeoffs [KingmaGao2023]. One can understand the noise schedule as defining a proposal distribution over the SNR (Figure 8), which determines the weight of a given noise level during training (Figure 7C–D). Hence, by applying an appropriate reweighting, one can transform samples from one schedule into those of another, since the weighting function converts the effective distribution over SNR levels [KingmaGao2023]. The most common schedule in SBI is the linear schedule from the original diffusion model formulation [HoJai2020a], which is defined as
@@
The cosine schedule [nichol2021improved] instead uses
@@
where $s$ is a small offset (typically $s=0$). This schedule smooths the allocation of noise across time steps, avoiding the imbalance of the linear schedule, and has become a widely adopted default for imaging tasks due to its empirical robustness in generative modeling.
Finally, the Elucidated Diffusion Model (EDM) schedule [karras2022elucidating], explicitly decouples training and sampling:
$\text{Training: }\lambda(t)=\mathcal{F}_{\mathcal{N}}^{-1}(1-t;2.4,2.4^{2})\quad\quad\text{Inference: }\lambda(t)=-2\rho\log\Bigl(\sigma_{\text{max}}^{1/\rho}+(1-t)\left(\sigma_{\text{min}}^{1/\rho}-\sigma_{\text{max}}^{1/\rho}\right)\Bigr),$ (49)
@@
A good schedule with high numerical accuracy requires the Jacobian of the velocity field to have a small Lipschitz norm [tsimpos2025optimal]. Empirically, high Lipschitz norms tend to occur near the boundaries $t\approx 0$ and $t\approx 1$. In practice, schedules are often truncated so that the SNR remains well-defined at the boundaries, using the rescaled time $\tilde{t}=t_{0}+(t_{1}-t_{0})t$. It is also common to condition on a normalized version of $\lambda_{t}$ instead of the raw time value $t$ for numerical stability, for example, by rescaling $\lambda_{t}$ to lie within [1,1][-1,1] [karras2022elucidating]. Recent work shows that, to reduce the variance as $t\to 0$, incorporating the tractable joint score $\nabla_{\boldsymbol{\theta}}\!\log p(\boldsymbol{\theta})$ in the objective function yields an unbiased estimator of the marginal score and improves simulation efficiency [de2024target, ko2025latent].
While most SBI papers still employ a linear schedule (LABEL:tab:papers), [Nautiyal2025-condisimConditionalDiffusionModelsSimulationBasedInferencef] showed that the cosine schedule outperforms the linear variant on the low-dimensional benchmark problems of [lueckmann2021benchmarking]. We additionally show that the EDM schedule outperforms the cosine schedule on the same problems (Section 6.1). Moreover, for neural likelihood estimation (NLE) of a PDE model, [haitsiukevich2024diffusion] demonstrated that the EDM schedule combined with an ODE sampler achieves superior performance compared to the linear schedule.

### v2

@@ 4.3.1 Multiscale Estimation
Many simulation-based inference problems involve high-dimensional, spatially structured parameters, where the resolution can be modeled hierarchically. So instead of multiple observations, we have one high-dimensional observation $\mathbf{y}_{\text{obs}}$, which can be mapped to different resolutions using a downscaling function $h(\mathbf{y}_{\text{obs}},r)$. We can then offload generation across multiple scales: we start at a coarse resolution and progressively upscale while refining detail. This idea was popularized in conditional image generation by cascaded diffusion pipelines that train separate (class-conditional) models at $(64\!\to\!128\!\to\!256)\,\text{px}$ resolution stages, conditioning each stage on the upscaled output of the previous one [72].
In the SBI setting, [29] combine diffusion models with wavelet decompositions to infer high-resolution velocity fields from seismic data. Here a diffusion model is learned for the top-level coarse component $p(\boldsymbol{\eta}^{(R)}\mid h(\mathbf{y}_{\text{obs}},R))$ and, for each resolution level $r$, a conditional detail model $p(\boldsymbol{\theta}^{(r)}\mid\boldsymbol{\eta}^{(r)},h(\mathbf{y}_{\text{obs}},r))$. This can be combined with an inverse wavelet transform to obtain the conditioning variable for the next finer stage $\operatorname{WT}^{-1}(\boldsymbol{\eta}^{(r)},\boldsymbol{\theta}^{(r)})=\boldsymbol{\eta}^{(r-1)}$. The joint posterior over all resolution levels factorizes as
$\begin{split}p(\{\boldsymbol{\eta}^{(r)}\}_{r=0}^{R},\{\boldsymbol{\theta}^{(r)}\}_{r=1}^{R}\mid\mathbf{y}_{\text{obs}})=\;&p(\boldsymbol{\eta}^{(R)}\mid h(\mathbf{y}_{\text{obs}},R))\;p(\boldsymbol{\theta}^{(R)}\mid\boldsymbol{\eta}^{(R)},h(\mathbf{y}_{\text{obs}},R))\\ &\times\prod_{r=1}^{R-1}p(\boldsymbol{\theta}^{(R-r)}\mid\boldsymbol{\eta}^{(R-r)},h(\mathbf{y}_{\text{obs}},R-r)),\end{split}$ (39)
so that training and ancestral sampling proceed naturally from coarse to fine scales. Each score estimator can be trained separately due to the exact factorization above and the analytic deconstruction path from high-resolution training data via $\operatorname{WT}(\boldsymbol{\eta}^{(r)})=(\boldsymbol{\eta}^{(r+1)},\boldsymbol{\theta}^{(r+1)})$. While the high-dimensional posterior has long-range spatial correlations, the individual conditional posteriors are closer to Gaussian distributions even when the high-dimensional posterior is not, making estimation of the individual factors easier [64]. This idea was first demonstrated with normalizing flows by [194], and [64] extended it to score-based generative models. Diffusion models are particularly well suited for this setting, as they scale to high dimensions and are very flexible in how the conditioning is used within the neural network architecture.
@@ 4.3.2 Causal Posterior Estimation
Causal posterior estimation extends simulation-based inference to settings where the posterior distribution inherits the structure of a causal model [40]. A key idea is to integrate the causal directed acyclic graph (DAG) directly into the network architecture of the diffusion model and the probability flow.
@@ 4.3.3 Joint Estimation
Instead of training only a conditional model for $p(\boldsymbol{\theta}\mid\mathbf{y})$, one can train a diffusion model on the joint $\mathbf{z}=(\boldsymbol{\theta},\mathbf{y})$ [58]. Using a transformer as the neural network architecture for the diffusion model, one can process any subset of factors forming the joint distribution. Yet, this architectural choice is prohibitive for constrained architectures (e.g., normalizing flows), as they typically require explicit conditioning on fixed inputs rather than being able to marginalize variables implicitly. In practice, the network is trained with random masking over components of $\mathbf{z}^{M_{C}}_{t}=(1-M_{C})\cdot\mathbf{z}_{t}+M_{C}\cdot\mathbf{z}_{0}$, where $M_{C}\in\{0,1\}^{d}$, leading to the modified objective
$\hat{s}=\operatorname*{argmin}_{s}\,\mathbb{E}_{M_{C}\sim p(M_{C}),\,\mathbf{z}_{0},\mathbf{z}_{t},t\sim p(\mathbf{z}_{0},\mathbf{z}_{t},t)}\left[\,\omega_{t}\,\|(1-M_{C})\cdot\bigl(s(\mathbf{z}^{M_{C}}_{t},t)-\nabla_{\mathbf{z}_{t}}\!\log p(\mathbf{z}_{t}\mid\mathbf{z}_{0})\bigr)\|_{2}^{2}\,\right].$ (42)
@@
Beyond posterior approximation, joint training also enables Bayesian model comparison using both the likelihood and the posterior [63]. Having access to all conditional distributions makes this approach particularly appealing for sequential neural score estimation. For example, [153] introduced SNPSE-C, which incorporates a score-based correction into the denoising score matching objective to refine posterior estimates iteratively. SNPSE-C requires access to likelihood and prior scores, which introduces additional simplifications and approximations, but these can be avoided when the diffusion model is trained directly on the joint. Extensions to spatial and non-parametric inference have also been proposed. For instance, [174] adapted the framework to spatially correlated data and demonstrated inference over multiple spatial conditional distributions simultaneously by providing temporal and spatial information as inputs to a transformer-based summary network.
@@ 4.3.4 Learning Scores Beyond Euclidean Spaces
The concept of the score function can be generalized beyond Euclidean parameter spaces to distributions defined on other structured domains. A detailed survey of diffusion models for data with special structures unrelated to SBI is provided by [191], covering cases such as discrete variables, symmetry-invariant data, and data constrained to manifolds. For targets with intrinsic geometric constraints, the score can be formulated directly on a manifold equipped with its metric structure. Examples include $\operatorname{E}(3)$-equivariant diffusion models [78], which respect rigid-body symmetries, and Riemannian score-based models [16], where both forward and reverse processes are defined with respect to the manifold geometry. [84] demonstrated this approach for the estimation of astrophysical orientation distributions on $\operatorname{SO}(3)$ extendable to neural likelihood estimation (NLE).
The score can also be defined for infinite-dimensional spaces when the target variable is a function rather than a finite-dimensional vector. [135] introduced diffusion models for Hilbert spaces and applied them to generative modeling of functions for Bayesian inverse problems and SBI. This formulation enables posterior estimation for high-dimensional objects like spatial fields, time series, or solutions to partial differential equations. A notable subtlety is that in infinite dimensions, the score matching objective is no longer equivalent to the standard denoising score matching loss, which requires careful reformulation of the training objective.
To address the fragmentation of generative objectives across different structured domains, Generator Matching reformulates the learning problem as matching the infinitesimal generator of a Markov process [75]. By targeting the linear operator governing the stochastic evolution rather than specific probability paths or noise distributions, this framework provides a unified objective for neural likelihood estimation on arbitrary state spaces, in particular it allows the construction of multi-modal generative models.
@@ 5 Design Choices for Diffusion Models in Simulation-Based Inference
An overarching decision concerns the inference backbone, namely the network responsible for sampling the target quantities $\mathbf{z}\sim p(\mathbf{z}\mid\mathbf{x})$. In principle, any architecture capable of handling the structure of the condition $\mathbf{x}$ is admissible, for instance, hybrid convolutional-recurrent models for time series [201], or permutation-invariant networks for sets [196, 99]. A common strategy for posterior estimation is to separate the model into a summary network and an inference network. The summary network compresses the data into a fixed-dimensional representation (i.e., an encoder or an embedding network), and the inference network is typically implemented as a stack of (residual) MLPs acting on this summary [139, 25]. However, this separation is optional; the flexibility of diffusion models allows these components to be folded into a single transformer architecture [58].
@@
Figure 7: Designing diffusion models for simulation-based inference (SBI). A Variance-preserving (VP) EDM schedule. B Variance-exploding (VE) EDM schedule (until $\sigma_{1}=131.9$). C Noise schedules of the log signal-to-noise ratio (SNR). The cosine and flow matching schedules exhibit similar relationships between time and log SNR, yet differ markedly in how the log SNR is weighted in the loss. D Flow matching emphasizes regions of low SNR more strongly than the other schedules. This can be adapted by sampling time $t$ from a power-law distribution with $\rho=-0.6$ rather than uniformly, which aligns the effective weighting with that of the EDM schedule. E Deterministic and stochastic latent trajectories from noise ($t=1$) to the posterior ($t=0$) for diffusion models trained with different noise schedules. The learned trajectories also reflect these design differences. Flow matching with optimal transport (OT) produces relatively straight paths of the probability flow from noise to the posterior, while EDM VP tends to give straighter paths than EDM VE. Stochastic sampling displays highly irregular motion. The example problem shown here is the conditional two moons [114] and the stochastic trajectories are smoothed with a simple moving average.
@@ 5.1 Variance of the Noise Schedule
A core design choice in diffusion models concerns the denoising process itself. The forward process (5) is characterized by the functions $\alpha_{t}$ and $\sigma_{t}$, which together specify the noise schedule. These functions determine how signal and noise are blended over time, and thereby shape both the difficulty of the denoising task and the model’s overall behavior [166]. In the variance-preserving (VP) parameterization,
@@
the total variance $\alpha_{t}^{2}+\sigma_{t}^{2}$ remains constant across $t$, assuming $\operatorname{Var}(\mathbf{z}_{0})=1$ and $0\leq\sigma_{t}\leq 1$ (Figure 7A). This ensures a balance between signal and noise at every step and underlies the original diffusion model formulation [71], which is still widely used in practice.
@@
so that the total variance grows throughout the forward process (Figure 7B). In some low-dimensional SBI inference tasks, VE has been reported to outperform VP [153, 58], while for high-dimensional imaging data, VP schedules generally perform better than VE schedules [166]. In our experiments Section 6.1, we find the VP schedule to be superior in most SBI tasks.
@@
which yields a lower total variance than the VP case for all $t$. For density estimation tasks, [166] observed that a sub-VP noise schedule can improve accuracy. However, most current SBI papers adopt either a VP schedule combined with a linear noise schedule or a VE schedule, and only a single paper employed a sub-VP schedule [198].
In practice, the choice of variance type interacts with both the noise schedule and the weighting function, jointly shaping the training difficulty over $t$, as we will discuss in the following sections. In our case studies, we find that for most problems, the VP schedule is a suitable choice (see Section 6). [113] show that the variance type does not fundamentally affect sample quality, since any formulation can be transformed into a VP schedule by rescaling $\alpha_{t}$ and $\sigma_{t}$ with $\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}$ and adjusting the weighting function accordingly. They introduce a trigonometric reparameterization, TrigFlow, which maps arbitrary schedules into VP form. Moreover, [171] proposed contractive variants of VP and sub-VP schedules that suppress the propagation of score and discretization errors, albeit at the cost of introducing possible bias.
@@ 5.2 Noise Schedule and Signal-to-Noise Ratio
The noise schedule determines how the log signal-to-noise ratio (SNR), $\lambda_{t}=\log(\alpha_{t}^{2}/\sigma_{t}^{2})$, evolves over time $t$, and thus controls the relative difficulty of denoising at different steps. Over the past few years, several schedules have been proposed, each with different motivations and practical tradeoffs [92]. One can understand the noise schedule as defining a proposal distribution over the SNR (Figure 8), which determines the weight of a given noise level during training (Figure 7C–D). Hence, by applying an appropriate reweighting, one can transform samples from one schedule into those of another, since the weighting function converts the effective distribution over SNR levels [92]. The most common schedule in SBI is the linear schedule from the original diffusion model formulation [71], which is defined as
@@
The cosine schedule [123] instead uses
@@
where $s$ is a small offset (typically $s=0$). This schedule smooths the allocation of noise across time steps, avoiding the imbalance of the linear schedule, and has become a widely adopted default for imaging tasks (with $s<\text{pixel bin size}$) due to its empirical robustness in generative modeling.
Finally, the Elucidated Diffusion Model (EDM) schedule [88], explicitly decouples training and sampling:
$\lambda_{\mathrm{train}}(t)$ $=\mathcal{F}_{\mathcal{N}}^{-1}(1-t;2.4,2.4^{2})$ (48)
$\lambda_{\mathrm{inference}}(t)$ $=-2\rho\log\Bigl(\sigma_{\text{max}}^{1/\rho}+(1-t)\left(\sigma_{\text{min}}^{1/\rho}-\sigma_{\text{max}}^{1/\rho}\right)\Bigr),$ (49)
@@
A good schedule with high numerical accuracy requires the Jacobian of the velocity field to have a small Lipschitz norm [177]. Empirically, high Lipschitz norms tend to occur near the boundaries $t\approx 0$ and $t\approx 1$. In practice, schedules are often truncated so that the SNR remains well-defined at the boundaries, using the rescaled time $\tilde{t}=t_{0}+(t_{1}-t_{0})t$. It is also common to condition on a normalized version of $\lambda_{t}$ instead of the raw time value $t$ for numerical stability, for example, by rescaling $\lambda_{t}$ to lie within [1,1][-1,1] [88]. Recent work shows that, to reduce the variance as $t\to 0$, incorporating the tractable joint score $\nabla_{\boldsymbol{\theta}}\!\log p(\boldsymbol{\theta})$ in the objective function yields an unbiased estimator of the marginal score and improves simulation efficiency [35, 96].
While most SBI papers still employ a linear schedule (LABEL:tab:papers), [121] showed that the cosine schedule outperforms the linear variant on the low-dimensional benchmark problems of [114]. We additionally show that the EDM schedule outperforms the cosine schedule on the same problems (Section 6.1). Moreover, for neural likelihood estimation (NLE) of a PDE model, [66] demonstrated that the EDM schedule combined with an ODE sampler achieves superior performance compared to the linear schedule.